# Distill-SynthKG: Distilling Knowledge Graph Synthesis Workflow for Improved Coverage and Efficiency

### v1

@@ Abstract
Knowledge graphs (KGs) generated by large language models (LLMs) are becoming increasingly valuable for Retrieval-Augmented Generation (RAG) applications that require knowledge-intensive reasoning. However, existing KG extraction methods predominantly rely on prompt-based approaches, which are inefficient for processing large-scale corpora. These approaches often suffer from information loss, particularly with long documents, due to the lack of specialized design for KG construction. Additionally, there is a gap in evaluation datasets and methodologies for ontology-free KG construction. To overcome these limitations, we propose SynthKG, a multi-step, document-level ontology-free KG synthesis workflow based on LLMs. By fine-tuning a smaller LLM on the synthesized document-KG pairs, we streamline the multi-step process into a single-step KG generation approach called Distill-SynthKG, substantially reducing the number of LLM inference calls. Furthermore, we re-purpose existing question-answering datasets to establish KG evaluation datasets and introduce new evaluation metrics. Using KGs produced by Distill-SynthKG, we also design a novel graph-based retrieval framework for RAG. Experimental results demonstrate that Distill-SynthKG not only surpasses all baseline models in KG quality—including models up to eight times larger—but also consistently excels in retrieval and question-answering tasks. Our proposed graph retrieval framework also outperforms all KG-retrieval methods across multiple benchmark datasets. We release the SynthKG dataset and Distill-SynthKG model publicly to support further research and development.
Distill-SynthKG: Distilling Knowledge Graph Synthesis Workflow for Improved Coverage and Efficiency
Prafulla Kumar Choubey∗1, Xin Su∗2, Man Luo∗2, Xiangyu Peng1 Caiming Xiong1, Tiep Le2, Shachar Rosenman2, Vasudev Lal 2 Phil Mui1, Ricky Ho1, Phillip Howard†2, Chien-Sheng Wu†1 1Salesforce Research, 2Intel Labs Correspondence: pchoubey@salesforce.com, xin.su@intel.com, man.luo@intel.com ∗Main authors contributed equally, †Senior authors
@@ 1 Introduction
Retrieval Augmented Generation (RAG) has gained widespread application for effectively connecting large language models (LLMs) with external knowledge sources. Recently, Knowledge Graph (KG) augmented RAG methods have demonstrated strong potential, offering several advantages such as effective corpus-level information summarization Edge et al. (2024), improved reasoning capabilities Gutiérrez et al. (2024); Li et al. (2024), and accurate modeling of historical customer issue resolutions for QA Xu et al. (2024).
@@
Recent works Edge et al. (2024); Gutiérrez et al. (2024) have begun exploring the use of LLMs to automate the construction of KGs, which then serve as knowledge sources for specific tasks such as question answering or building intelligent agentic frameworks. However, these existing approaches have several limitations. First, they rely on simple zero-shot or few-shot in-context learning methods to construct knowledge graphs in a single step using LLMs like GPT-4o OpenAI (2024). Consequently, such approaches can incur significant inference costs when applied across large corpora due to the need for many commercial API calls. These methods also lack a rigorous and reliable design specifically tailored for KG construction. Having LLMs process entire documents, particularly long texts, has been shown to potentially lead to issues such as information loss Edge et al. (2024). Second, there is a lack of existing datasets or evaluation methods to effectively evaluate document-level ontology-free KGs. This absence makes it difficult to identify whether errors in RAG systems stem from issues in specific reasoning components or from poor-quality KGs that propagate errors throughout the system.
@@
To address these limitations, we introduce SynthKG, a novel LLM-based KG construction workflow. We further distill this workflow into a smaller LLM named Distill-SynthKG, which enables efficient, one-step generation of high-quality document-level KGs. In SynthKG, we begin by splitting the input document into manageable, semantically complete text chunks. Each chunk is then processed through a decontextualization step where entity disambiguation occurs based on the previous context, making each chunk an independent, self-contained unit. We then prompt the LLM to extract entities, relations, and relevant propositions from each text chunk, which are combined to form the final KG. Finally, we fine-tune our smaller Distill-SynthKG LLM on the KGs produced by SynthKG, enabling it to generate the KG for a given document in a single inference step.
Additionally, we propose a method for constructing an evaluation dataset for document-level ontology-free KGs, along with a corresponding KG evaluation framework. Specifically, we re-purpose existing multihop QA datasets by converting questions and answers into ground truth relation triplets, where the answer appears as either the head, tail, or predicate in a triplet. Using these ground truth triplets for each document, we introduce semantic similarity and keyword-based metrics to assess the coverage of triplets from a KG.
@@
Finally, we present a new graph-based retrieval framework based on the KGs generated by Distill-SynthKG. We design a progressive retrieval method that begins with proposition retrieval, leveraging the graph structure to retrieve related triplets, propositions, and text chunks relevant to the input query. Our proposed retriever outperforms state-of-the-art retrieval methods in both retrieval accuracy and question-answering accuracy, showing improvements across three multihop QA datasets: MuSiQue Trivedi et al. (2022), 2WikiMultiHopQA Ho et al. (2020), and HotpotQA Yang et al. (2018). Furthermore, our KG coverage evaluation framework correlates strongly with both QA and retrieval performance, demonstrating its effectiveness in evaluating document-level KG coverage.
In summary, our main contributions are as follows:
@@
We introduce SynthKG, a novel LLM-based workflow that generates high-quality, high-coverage document-level ontology-free KGs.
We introduce Distill-SynthKG, which leverages SynthKG to synthesize training data and fine-tune a much smaller LLM. This approach simplifies the multi-step process into a single-step inference, significantly reducing the reliance on LLM calls.
@@
We propose new KG evaluation datasets by re-purposing existing multi-hop QA datasets and introducing new evaluation metrics.
@@
We introduce a novel graph-based retrieval method that leverages KGs generated by Distill-SynthKG.
@@
Our experiments across multiple datasets demonstrate that Distill-SynthKG not only produces KGs of higher quality than all baselines—including models up to eight times larger—but also consistently outperforms them in retrieval and question-answering tasks. Furthermore, the proposed graph-based retrieval framework surpasses all baseline KG-based retrieval methods.
@@ 2 Related Work
Recently, there has been a growing interest in using Knowledge Graphs (KG) for different Retrieval-Augmented Generation (RAG) applications. For instance, GraphRAG Edge et al. (2024) shows the advantages of using KGs over a text corpus for answering global queries that require summarizing information from multiple documents. HippoRAG Gutiérrez et al. (2024) demonstrates that applying personalized PageRank algorithms on LLM-derived KG can enhance retrieval accuracy for complex multi-hop reasoning questions. GraphReader Li et al. (2024) shows how KGs can enable LLM agents to plan and reason in a long context to answer complex questions. These approaches focus on maximizing KG utility.
All the above work, along with many others such as Chia et al. (2022); Trajanoska et al. (2023); Chen and Bertozzi (2023); Kai Zhang (2023); Nayak and Timmapathini (2023); Mihindukulasooriya et al. (2023); Zhu et al. (2024); Jiao et al. (2023); Khorashadizadeh et al. (2023); Han et al. (2024); Yao et al. (2024); Bi et al. (2024); Ding et al. (2024); Sanmartin (2024); Sun et al. (2024); Yao et al. (2023); Chase (2022) have used LLM prompting to build KGs or extract semantic relation triplets from text. However, all prior works have overlooked improving the efficiency of ontology-free KG construction. We are the first to develop a specialized LLM for KG construction, enhancing efficiency by shifting from large, advanced LLMs to smaller, more efficient models without sacrificing performance.
@@ 3 Distill-SynthKG
We present Distill-SynthKG, a framework for fine-tuning LLMs by distilling the multi-step KG synthesis process (SynthKG) into a streamlined, single-step approach. This allows for the direct generation of KGs from documents using smaller-scale LLMs. Specifically, we first apply SynthKG to generate KGs from documents using a larger LLM. We then distill this process by training a smaller LLM on the resulting document-KG pairs, producing the distilled model, Distill-SynthKG.
3.1 SynthKG
@@
SynthKG consists of two main steps: (1) document chunking and decontextualization, followed by (2) entity, relation and proposition extraction. These steps ensure high coverage of extracted entities and relations while minimizing information loss. We present an overview of SynthKG in Figure 1. Additionally, we provide details of all LLM prompts utilized in this process in Appendix A.
3.1.1 Document Chunking and Decontextualization
@@
Directly inputting long texts into an LLM has been shown to result in information loss Edge et al. (2024). To mitigate this risks, we first split each input document into smaller, more manageable chunks before processing them with the LLM in subsequent steps. This chunking is done along sentence boundaries, without overlap, to preserve semantic coherence and avoid redundancy.
@@
However, processing each chunk in isolation can lead to a loss of prior context. For example, if "John Doe" appears in one chunk and "John" in another, we might lose track of who "John" refers to. To prevent this, we apply a "decontextualization" step, where we prompt the LLM to rewrite each chunk, replacing all entity mentions with their most informative form based on the context of the preceding chunk. For example, if “John Doe” is introduced in a previous chunk, subsequent mentions of “John D.” “John,” or related pronouns are replaced with “John Doe.” This approach not only preserves context but also prevents the same entity from being represented in different forms, which could lead to redundancy, discontinuous KG paths, and reduced accuracy during KG inference or retrieval. The first chunk of a document is not decontextualized, as chunking does not lead to context loss in this case. We provide an example of decontextualized chunk in Figure 9.
One potential downside of prompting the LLM to rewrite chunks is that the rewritten version may deviate from the original, potentially introducing information loss or hallucination. To address this, we use ROUGE scores to compare the original and decontextualized chunks and filter out those that show significant deviations due to the decontextualization process. Detailed experimental settings are provided in the Section 6.1.
3.1.2 Entity and Relation Extraction
@@
Following the paradigm of Edge et al. (2024) and Gutiérrez et al. (2024), we first prompt the LLM to extract all entities and their corresponding types from each text chunk, as shown in Step 3 of Figure 1. Then, we prompt the LLM again to generate all propositions and corresponding relation triplets based on the text chunk and previously extracted entities. Each relation is represented by quadruplets consisting of a source entity, predicate, target entity, and a proposition (see Figure 1 for examples). The proposition is a sentence that describes the semantic relation between the source and target entities, encapsulating all key details of that relation.
@@
We extend traditional KG triples by adding a proposition component. One reason for this is that propositions function as an intermediate chain of thought Wei et al. (2022), enabling the LLM to first articulate the relevant context coherently before extracting the corresponding triplets. This approach therefore better leverages contextual information. Additionally, the proposition acts as a fine-grained, self-contained retrieval unit, which facilitates the construction of KG-based retrieval indices. Beyond triplets and text chunks, our final KG incorporates these clear, independent propositions. For example, the proposition “OWC Pharmaceutical Research Corp preferred stock is convertible to common stock at $0.20 per share.” provides important contextual details, such as the “conversion price $0.20 per share,” and also serves as a precise, indexable unit.
@@ 3.2 Distilling SynthKG
While the detailed, chunk-by-chunk approach in SynthKG enables the generation of high-quality KGs using LLMs, it introduces efficiency challenges. Each time a KG is constructed from a document, we must repeatedly prompt an LLM, resulting in substantial computational or API costs, which hinders the scalability of KG construction. For example, processing a 1000-word document involves 12 LLM inference calls (4 chunks, with 3 calls per chunk for decontextualization, entity extraction, and relation extraction). To mitigate this, we distill SynthKG into a smaller LLM by leveraging the document-KG pairs produced in the original process. Specifically, we fine-tune a smaller LLM to take entire documents as input, enabling it to directly generate the same KGs as SynthKG (i.e., a series of quadruplets) in a single inference step.
We hypothesize that the high-quality document-KG pairs generated by SynthKG can be effectively used to train smaller LLMs, helping to mitigate the information loss that commonly occurs when processing entire documents without such training. However, there is currently no large-scale dataset available for this type of training, making SynthKG essential for creating the data necessary to enable such model distillation.
@@ 4 KG Coverage Evaluation
Evaluating the quality of the extracted KG is essential, as it directly impacts its downstream applications. However, there is a lack of document-level data for KG evaluation. Although DocRED Yao et al. (2019) is one existing dataset, it is limited to just 96 relations, making it less suitable for KGs used in retrieval tasks, which often rely on an open ontology and include diverse relations. To address this gap, we propose generating proxy ground truth relation triplets from multihop QA datasets and introduce metrics for evaluating the coverage of these proxy triplets in the extracted KG.
@@
We use GPT-4o to generate triplets from QA pairs. Given that multihop QA requires reasoning over multiple facts, we generate one triplet for each individual fact. In datasets where these facts are present as subquestion-answer pairs, we create triplets using these pairs while ensuring that the answer is used as the head, relation, or tail in the triplet. In cases where facts or subquestions are unavailable, we use GPT-4o to first generate the required subquestions before subsequently generating the corresponding triplets. The prompts used for generating the triplets and decomposed questions, along with relevant examples, are provided in Appendix B.1. In human evaluations, we found a high degree of validity in triplets extracted by GPT-4o with this approach; see Appendix B.2 for details.
@@
Existing KG evaluation metrics typically depend on exact match or F1 score at the text level, given that relations are derived from a predefined set. However, this approach is inefficient for ontology-free KGs, where entities and relations are not constrained. To address this, we use semantic matching to align the extracted triplets with the ground truth triplets, and propose three complementary metrics: semantic scores, triplet coverage, and F1 scores. Note that the quality of an extracted KG is evaluated based on its coverage of a given set of ground truth triplets. Therefore, these three metrics are not aimed at measuring the graph’s comprehensiveness but rather at verifying whether the important triplets—those critical for answering questions—are included in the graph. As a proxy for comprehensiveness, we additionally compare the number of extracted triplets. Our three proposed metrics are defined as follows:
Semantic score: We calculate the cosine similarity between the vector representation of each ground truth triplet and the triplets in the knowledge graph, taking the highest similarity score as the semantic score for that ground truth triplet. A higher semantic score indicates a closer match between the ground truth and the extracted graph.
Triplet Coverage: We use a cutoff threshold: if the semantic score for a ground truth triplet exceeds this threshold, it is marked as covered (coverage = 1); otherwise, the triplet is not covered (coverage = 0).
F1 score: We use the semantic score to identify the triplet from the knowledge graph that most closely matches the ground truth triplet. Then, we compute the F1 score by comparing the text of the extracted triplet with that of the ground truth triplet.

### v2

# Scaling Knowledge Graph Construction through Synthetic Data Generation and Distillation

@@ Abstract
Document-level knowledge graph (KG) construction faces a fundamental scaling challenge: existing methods either rely on expensive large language models (LLMs), making them economically nonviable for large-scale corpora, or employ smaller models that produce incomplete and inconsistent graphs. We find that this limitation stems not from model capabilities but from insufficient training on high-quality document-level KG data. To address this gap, we introduce SynthKG, a multi-step data synthesis pipeline that generates high-quality document-KG pairs through systematic chunking, decontextualization, and structured extraction using LLMs. By fine-tuning a smaller LLM on synthesized document-KG pairs, we streamline the multi-step process into a single-step KG generation approach called Distill-SynthKG. Furthermore, we repurpose existing question-answering datasets to construct KG evaluation datasets and introduce new evaluation metrics. Using KGs produced by Distill-SynthKG, we also design a novel graph-based retrieval framework for RAG. Experimental results demonstrate that Distill-SynthKG not only surpasses all baseline models in KG quality (including models up to eight times larger) but also consistently improves in retrieval and question-answering tasks. Additionally, our proposed graph retrieval framework outperforms all KG-retrieval methods across multiple benchmark datasets.
@@ 1 Introduction
Retrieval Augmented Generation (RAG) has been widely adopted for effectively connecting large language models (LLMs) with external knowledge sources. Recently, Knowledge Graph (KG)-augmented RAG methods have demonstrated strong potential, offering several advantages such as effective corpus-level information summarization (Edge et al., 2024), improved reasoning capabilities (gutiérrez2024hipporag; Li et al., 2024), and accurate modeling of historical customer issue resolutions for QA (Xu et al., 2024).
@@
Recent works (Edge et al., 2024; gutiérrez2024hipporag) have begun exploring the use of LLMs to automate the construction of KGs, which then serve as knowledge sources for tasks such as question answering or building intelligent agentic frameworks. However, these approaches face a fundamental scaling challenge: they rely on simple zero-shot or few-shot prompting to construct KGs in a single step using LLMs such as GPT-4o (OpenAI, 2024). Consequently, they can incur significant inference costs when applied to large corpora due to the need for many commercial API calls. These methods also lack a rigorous, reliable design tailored to document-level KG construction, and processing entire documents with LLMs, particularly for long texts, can lead to information loss (Edge et al., 2024). Finally, the lack of datasets and evaluation methods for document-level, ontology-free KGs makes it difficult to determine whether errors in RAG systems stem from failures in specific reasoning components or from low-quality KGs that propagate errors throughout the system.
@@
We find that these limitations stem not from model capabilities but from the absence of high-quality training data for document-level KG construction. While other structured extraction tasks benefit from supervised datasets, ontology-free KG construction lacks substantial training corpora, forcing reliance on expensive zero-shot inference. To address this data gap, we introduce SynthKG, a data synthesis pipeline for KG construction. We further distill this pipeline into a smaller LLM, Distill-SynthKG, which enables efficient one-step generation of high-quality document-level KGs. In SynthKG, we begin by splitting the input document into manageable, semantically complete text chunks. Each chunk is then processed through a decontextualization step that performs entity disambiguation using prior context, making each chunk a self-contained unit. We then prompt the LLM to extract entities, relations, and relevant propositions from each chunk, which are combined to form the final KG. By fine-tuning Distill-SynthKG on the synthetic document–KG pairs produced by SynthKG, smaller models can generate high-quality KGs for a given document in a single inference step.
@@
Additionally, we propose a method for constructing evaluation datasets for document-level, ontology-free KGs, along with a corresponding evaluation framework. Specifically, we repurpose existing multihop QA datasets by converting questions and answers into ground-truth relation triplets, where the answer appears as the head, tail, or predicate of a triplet. Using these ground-truth triplets for each document, we introduce semantic-similarity and keyword-based metrics to assess KG triplet coverage. Finally, we present a graph-based retrieval framework built on KGs generated by Distill-SynthKG. We design a progressive retrieval method that begins with proposition retrieval and leverages the graph structure to retrieve related triplets, propositions, and text chunks relevant to the query. Our retriever outperforms state-of-the-art methods in both retrieval and question-answering accuracy across three multihop QA datasets: MuSiQue (Trivedi et al., 2022), 2WikiMultiHopQA (Ho et al., 2020), and HotpotQA (Yang et al., 2018). Moreover, our KG coverage evaluation framework correlates strongly with both QA and retrieval performance, demonstrating its effectiveness for evaluating document-level KG coverage.
In summary, our contributions are as follows:
1.
@@
We introduce SynthKG, a systematic data synthesis pipeline that generates high-quality document-level ontology-free KGs, addressing the training data scarcity in KG construction.
2.
We present Distill-SynthKG, demonstrating that smaller LLMs can achieve large-model KG construction quality when trained on appropriate synthetic data.
3.
@@
We establish comprehensive KG evaluation datasets and metrics by re-purposing existing multi-hop QA datasets.
4.
@@
We design a novel graph-based retrieval framework that effectively leverages KG structure for RAG.
5.
@@
Our experiments show that Distill-SynthKG produces higher-quality KGs than all baselines, including models up to eight times larger, while consistently improving retrieval and question-answering performance.
@@ 2 Related Work
Recently, there has been a growing interest in using KGs for different Retrieval-Augmented Generation (RAG) applications. For instance, GraphRAG (Edge et al., 2024) shows the advantages of using KGs over a text corpus for answering global queries that require summarizing information from multiple documents. HippoRAG (gutiérrez2024hipporag) demonstrates that applying personalized PageRank algorithms on LLM-derived KG can enhance retrieval accuracy for complex multi-hop reasoning questions. GraphReader (Li et al., 2024) shows how KGs can enable LLM agents to plan and reason in a long context to answer complex questions. These approaches focus on maximizing KG utility.
All the above work, along with many others such as Chia et al. (2022); Trajanoska et al. (2023); Chen and Bertozzi (2023); Kai Zhang (2023); Nayak and Timmapathini (2023); Mihindukulasooriya et al. (2023); Zhu et al. (2024); Jiao et al. (2023); Khorashadizadeh et al. (2023); Han et al. (2024); Yao et al. (2024); Bi et al. (2024); Ding et al. (2024); Sanmartin (2024); Sun et al. (2024); Yao et al. (2023); Chase (2022) have used LLM prompting to build KGs or extract semantic relation triplets from text. However, all prior works have overlooked improving the efficiency of ontology-free KG construction. We are the first to develop a specialized LLM for KG construction, enhancing efficiency by shifting from large models to smaller, more efficient models without sacrificing performance.
@@ 3 Distill-SynthKG
We present Distill-SynthKG, a data-centric approach for scaling KG construction. Rather than relying on increasingly larger models, we generate high-quality training data through a systematic pipeline (SynthKG) and use it to train smaller models for document-level KG construction. This enables efficient single-step KG generation from documents using smaller-scale LLMs while matching the quality of expensive multi-step pipelines.
3.1 SynthKG: A Data Synthesis Engine
@@
Traditional prompt-based KG extraction treats each document as an isolated zero-shot problem, leading to inconsistent outputs and preventing systematic learning. SynthKG addresses this by decomposing KG construction into reproducible stages that generate consistent, high-quality training data. The pipeline consists of two main steps: (1) document chunking and decontextualization, followed by (2) entity, relation and proposition extraction. These steps ensure high coverage of extracted entities and relations while minimizing information loss. We present an overview of SynthKG in Figure˜1, with detailed prompts in Appendix˜C.
Document Chunking and Decontextualization
@@
Directly inputting long texts into an LLM has been shown to result in information loss (Edge et al., 2024). To mitigate this risk, we first split each input document into smaller, more manageable chunks before processing them with the LLM in subsequent steps. This chunking is done along sentence boundaries, without overlap, to preserve semantic coherence and avoid redundancy.
@@
However, processing each chunk in isolation can lead to a loss of prior context. For example, if “John Doe” appears in one chunk and “John” in another, we might lose track of who “John” refers to. To prevent this, we apply a “decontextualization” step, where we prompt the LLM to rewrite each chunk, replacing all entity mentions with their most informative form based on the context of the preceding chunk. This step serves a critical dual purpose: ensuring entity consistency across chunks and creating self-contained text units that can be independently processed. For example, if “John Doe” is introduced in a previous chunk, subsequent mentions of “John D.” “John,” or related pronouns are replaced with “John Doe.” This not only preserves context but also prevents the same entity from being represented in different forms, which could lead to redundancy, discontinuous KG paths, and reduced accuracy at inference time. The first chunk of a document is not decontextualized, as chunking does not lead to context loss in this case. We provide an example of a decontextualized chunk in Figure˜8.
To verify that the preceding chunk is sufficient for decontextualization, we calculate the average chunk distance for the same entity within each document in our generated dataset of 100K samples (details of this dataset are described in Section˜6.1). Specifically, we measure the distance between the first occurrence of each entity and its subsequent mentions. The overall average chunk distance per entity is 0.9, indicating that, on average, entities are mentioned again within less than one chunk after their first mention. This suggests that using only the preceding chunk is sufficient and that no significant number of entities remain unresolved due to the chunk-based decontextualization process.
A potential drawback of prompting an LLM to rewrite chunks is that the rewritten text can drift from the source, introducing information loss or hallucinations. To mitigate this risk, we compute ROUGE scores between the original and decontextualized chunks and filter out cases with substantial divergence. Full experimental settings are reported in Section˜6.1.
To further evaluate decontextualization accuracy, we manually annotate 75 randomly selected rewritten chunks. Three authors each label 25 chunks, with access to both the original chunk and the full document. For each chunk, annotators assess whether any modifications were made, whether those modifications are correct, and whether any information was lost.
Across the 75 chunks, we record 593 edits in total, of which only six involve incorrect modifications and four exhibit information loss. Overall, these results suggest that decontextualization produces high-quality, self-contained text. Qualitatively, most edits increase specificity, for example, replacing a generic term like “scientists” with “Darwinian scientists,” indicating that the rewritten chunks are typically understandable without additional context.
Entity and Relation Extraction
@@
Similar to Edge et al. (2024) and gutiérrez2024hipporag, we first prompt the LLM to extract all entities and their corresponding types from each text chunk, as shown in Step 3 of Figure˜1. Then, we prompt the LLM again to generate all propositions and corresponding relation triplets based on the text chunk and previously extracted entities. Each relation is represented by quadruplets consisting of a source entity, predicate, target entity, and a proposition (see Figure˜1 for examples). The proposition is a sentence that describes the semantic relation between the source and target entities, encapsulating all key details of that relation.
@@
We extend traditional KG triples by adding a proposition component, which functions as an intermediate chain of thought (Wei et al., 2022) enabling the LLM to first articulate the relevant context coherently before extracting the corresponding triplets. This approach therefore better leverages contextual information. Additionally, the proposition acts as a fine-grained, self-contained retrieval unit, which facilitates the construction of KG-based retrieval indices. Beyond triplets and text chunks, our final KG incorporates these clear, independent propositions. For example, the proposition “OWC Pharmaceutical Research Corp preferred stock is convertible to common stock at $0.20 per share.” provides important contextual details, such as the “conversion price $0.20 per share,” and also serves as a precise, indexable unit.
Crucially, this multi-step decomposition creates consistent patterns that can be learned. Unlike single-step prompting which produces variable outputs, our pipeline generates KGs with predictable structure: consistent entity naming from decontextualization, systematic coverage from chunk-by-chunk processing, and interpretable intermediate representations through propositions. These properties make the outputs suitable as training data for smaller models.
@@ 3.2 Distilling SynthKG
While SynthKG’s detailed, chunk-by-chunk workflow enables high-quality KG construction with LLMs, it creates efficiency challenges. Building a KG from a single document requires multiple LLM calls, which increases computation or API cost and limits scalability. For instance, processing a 1,000-word document requires 12 inference calls: the document is split into four chunks, and each chunk triggers three calls for decontextualization, entity extraction, and relation extraction.
To scale KG construction, we distill the multi-step SynthKG pipeline into a single-step model, as shown in Figure˜1. The key insight is that once KG construction is decomposed into systematic stages, it can be learned as a pattern recognition problem rather than re-solved through multi-step prompting at inference time. Concretely, we fine-tune a smaller LLM on the document–KG pairs generated by SynthKG, enabling it to process full documents end-to-end and produce high-quality KGs in a single forward pass.
This distillation approach overcomes the limitations of direct prompting because the synthetic training data provides two essential signals: (1) consistent examples of handling long documents without losing information, learned from the chunk-aggregated outputs; and (2) an implicit encoding of the multi-step procedure into the model parameters. As a result, the model learns not only to extract triplets, but also to preserve entity consistency and the coverage patterns demonstrated by SynthKG. Notably, no large-scale dataset currently exists for this form of document-to-KG training, making SynthKG crucial for generating the data required to enable effective distillation.
@@ 4 KG Coverage Evaluation
Evaluating the quality of the extracted KG is essential for our data-centric approach, as it enables systematic assessment of both the synthetic training data and the distilled models. However, there is a lack of document-level evaluation resources for ontology-free KGs. Although DocRED (Yao et al., 2019) is one existing dataset, it is limited to just 96 relations, making it unsuitable for open-domain KGs that require diverse, unconstrained relations. To address this gap and enable scalable evaluation, we propose a framework that re-purposes multihop QA datasets to create proxy ground truth triplets, providing the first systematic evaluation method for document-level ontology-free KG construction.
@@
We leverage multihop QA datasets because they naturally encode the structured facts needed to answer complex questions, which is exactly the kind of information KGs should capture. We use GPT-4o to generate triplets from QA pairs, since each multihop question implicitly contains multiple interconnected facts. When a dataset provides these facts as subquestion–answer pairs, we form triplets directly from those pairs and ensure that the final answer appears as the head, relation, or tail of at least one triplet. When subquestions or supporting facts are not available, we first prompt GPT-4o to generate the necessary subquestions and then generate the corresponding triplets.
Although this procedure is synthetic, it yields a scalable and consistent evaluation framework that would be impractical to obtain through manual annotation at the scale required for training data validation. The prompts for triplet generation and question decomposition, along with illustrative examples, are provided in Appendix D.1. Human evaluation shows that the resulting triplets are largely valid (86% accuracy); see Appendix D.2 for details.
@@
Existing KG evaluation metrics typically rely on exact match or token-level F1, often assuming relations come from a predefined schema. This assumption breaks down for ontology-free KGs, where entity and relation strings are unconstrained and surface-form variation is common. To handle this setting, we perform semantic matching to align extracted triplets with ground-truth triplets and report three complementary measures: semantic scores, triplet coverage, and F1.
We emphasize coverage over precision for a practical reason. In RAG applications, missing critical information (low recall) is usually more harmful than extracting additional facts (lower precision): irrelevant triplets can often be ignored during retrieval, whereas absent triplets can make relevant questions unanswerable. Accordingly, our metrics focus on whether the graph includes the key triplets needed to answer questions. As a proxy for overall comprehensiveness, we also report the total number of extracted triplets.
Our three proposed metrics are defined as follows:
Semantic score: We calculate the cosine similarity between the vector representation of each ground truth triplet and the triplets in the KG, taking the highest similarity score as the semantic score for that ground truth triplet. A higher semantic score indicates a closer match between the ground truth and the extracted graph.
Triplet Coverage: If the semantic score for a ground truth triplet exceeds a cutoff threshold, it is marked as covered (coverage = 1); otherwise, the triplet is not covered (coverage = 0).
F1 score: We use the semantic score to identify the triplet from the KG that most closely matches the ground truth triplet. Then, we compute the F1 score by comparing the text of the extracted and ground truth triplets.
These metrics collectively provide a comprehensive view of KG quality: semantic score captures conceptual alignment, coverage measures recall of critical information, and F1 score assesses surface-level accuracy. Together, they enable systematic comparison across different KG construction methods and validate the effectiveness of our data-centric approach.